# Accommodating Fabrication Defects on Floquet Codes with Minimal Hardware Requirements

### v1

@@ 1 Introduction
Floquet codes are a relatively new and exciting area in quantum error correction (QEC). From a theoretical perspective, they have a mathematical structure which is distinct from other families of codes such as stabiliser and subsystem codes, which lends them some interesting and unique properties [1]. In addition, from a more practical perspective, Floquet codes benefit from considerably reduced connectivity requirements in hardware, compared to the surface or colour codes, while maintaining a reasonably high threshold under Pauli noise and a low tera-quop footprint for certain noise models [2, 3, 4].
@@
We numerically benchmark the performance of our method by implementing circuits for a planar honeycomb code memory experiment (specifically considering the model of Gidney et al. [3]) in Stim [25], a fast Clifford circuit simulator, and correcting for Pauli errors on the adapted code using PyMatching, a matching decoder [26]. We find considerable robustness to fabrication defects: we determine a defect threshold of approximately $13\%$, and numerically demonstrate that Pauli error thresholds exist for defect probabilities close to this threshold. Hence, the region of parameter space in which the code remains fault-tolerant is large and comparable to that of the surface code [17]. The reduction in Pauli error threshold with defect probability is also quantitatively similar to previous findings for the surface code [19]. This is quite surprising, as the honeycomb code achieves this with more restrictive constraints on the connectivity of the device.
We further introduce two methods which take advantage of the underlying hardware. First, we propose even better methods for accommodating defects in hardware in which extra connectivity is present, such as on square lattices [27, 28, 29]. And secondly, we propose a method for additionally accommodating defective auxiliary qubits in lattices in which those are present—or, more generally, defective connections between data qubits—such as the heavy hex lattice [30].
The rest of this paper is laid out as follows. In Sec. 2 we provide some background on the honeycomb code and fabrication defects. In Sec. 3 we describe our strategy for removing defective qubits from the code lattice. Results from simulations are described in Sec. 4. We propose some improvements that can be made depending on the hardware of the quantum processor in Sec. 5. Finally, we conclude with some future research directions in Sec. 6.
@@ 2 Background
A Floquet code is defined by a series of check measurements and a sequence in which they are measured [^1] For our purposes, we do not distinguish between finite or infinite-period dynamical codes, as other authors do [31, 32]; they will both be, in principle, amenable to our techniques.. Many Floquet codes are definable on the lattice of a colour code [23]. We focus here on the honeycomb code [1, 2], however this example can be easily generalised.
@@ 2.1 Honeycomb code
The honeycomb code is a Floquet code defined on a honeycomb lattice [^2] This can be generalised to any lattice that is 3-face-colourable and tri-valent, i.e., where each vertex is adjacent to three edges.. Each vertex represents a physical qubit, and each edge represents a two-qubit Pauli measurement between neighbouring qubits. An example lattice is presented in Fig. 1, with the plaquettes coloured red, blue and green in such a way that no neighbouring plaquettes share the same colour. Edges are also assigned colours according to the plaquettes they connect; for example a red edge connects a pair of red plaquettes. A key benefit that this code has over other error-correcting codes, such as the surface code, is the low connectivity requirements of the code: each qubit only needs to be connected to three nearest neighbours, and each measurement is only between pairs of adjacent qubits. This makes Floquet codes appealing for architectures with low connectivity, such as IBM’s heavy-hex lattice [30], as well as devices with native two-body measurements such as Majorana devices [33, 4].
@@
The single-qubit measurements anti-commute with plaquettes along the boundary. The result of this is that some stabilisers are non-deterministic at times and therefore cannot form detectors. However, this does not seem to significantly affect the code’s performance when compared to the periodic honeycomb code [3].
We form detectors from two measurements of plaquette operators. For example, after the sequence $B\to R\to B$, we measure all bulk green plaquettes twice (once during $B\to R$, and a second time during $R\to B$), therefore detectors can be formed for each green plaquette by comparing the two measured values. Measurement results from detectors are deterministic in the presence of zero noise, meaning that a decoder can use these results to identify and correct errors when noise is present. Boundary plaquettes are measured non-deterministically when the single-qubit edge measurements precede the two-qubit edge measurements. However, they are re-measured deterministically (assuming no error has occurred) when the two-qubit edge measurements precede the single-qubit measurements. For example, a green boundary plaquette in Fig. 1 is first measured non-deterministically when the blue edge measurements immediately precede the red edge measurements ($B\to R$). It is then measured deterministically for a second time after the sequence $R\to B$, at which point we can form a detector comparing these two measurement outcomes.
@@ 2.2 Fabrication defects
Fabrication defects have been studied extensively in the context of surface codes, and can be accommodated by merging adjacent stabilisers together into “super-stabilisers” in order to remove the defective qubit(s) [17, 18, 19, 20]. These super-stabilisers are measured by introducing “gauge operators” around a defect, such that the super-stabiliser measurements can be formed by taking products of gauge measurements. This is similar to how edge measurements combine to form plaquette measurements in the honeycomb code, with the difference being that there is no dependency on the order in which gauge measurements are performed. However, translating this procedure to the honeycomb code is non-trivial; while each sub-round of the honeycomb code can be mapped to an instance of a surface code [1], the surface codes evolve with each sub-round.
Aasen et al. have investigated ways of accommodating fabrication defects on honeycomb and 4.8.8 Floquet codes [22]. They propose several strategies, two of which involve removing all qubits in a plaquette adjacent to a defect. An important consideration for such methods is whether the device has extra connectivity available for the formation of new plaquettes, which is required for some methods proposed [22]. Respecting the constraints of hardware while removing the smallest number of qubits and edges from the lattice possible is important for maintaining a high percolation threshold, and presumably a high Pauli noise threshold under finite fabrication defect rates***One of the methods proposed in Ref. [22], while requiring extra connectivity, only necessitates the removal of a single qubit, which could be useful for honeycomb code implementations. However, the properties of the resulting graph and the updated measurement schedule make this strategy potentially hard to generalise to other Floquet codes (including the honeycomb code with boundaries) or to multiple defects in the lattice..
@@ 3 Strategy for removing defective qubits from the lattice
Here we outline the strategy for removing defective qubits from the lattice. The main idea is to create “super-plaquettes” which avoid using the defective qubits, thereby removing them from the lattice. We do this by merging some plaquettes in the original lattice together, while other existing plaquettes are replaced by two-body plaquettes. We will do all this while retaining the 3-face-colourability and tri-valence of the graph, for ease of generalisation to other Floquet codes. To achieve this, we need to remove a pair of qubits along an edge, as described below. There is some subtlety around choosing the extra removed qubit (or qubits if there is more than one defect) in such a way that maximises the distance of the resulting code—we shall refer to this as the “edge choice”, since choosing a qubit to remove is the same as choosing an edge adjacent to the given fabrication defect. We also do not allow for extra connectivity between qubits in the updated Floquet code. In Appendix A, we describe how improved results can be obtained in the presence of extra connections in the underlying hardware.
@@ 3.1 The super-plaquette algorithm
The final result of this process is shown in Fig. 2(c). Note that the defective qubit and its neighbour are no longer connected to any other qubits, nor do they contribute to any plaquettes in the code, therefore they can now be removed.
@@
The resulting super-plaquette is of the same type as the merge plaquettes: it is formed from a cycle of edges of alternating type not equal to $D$ (alternating $YY$ and $ZZ$ measurements for the example in Fig. 2). This follows from the fact that the cycle of edges that wraps around the super-plaquette is formed of edges from the set $C$ and edges from merge plaquettes, none of which have basis equal to the merge plaquette type. To see that the cycle of bases is alternating, consider the edges adjacent to $(E,B)\in C$ in Fig. 2(b). There are five edges adjacent to $E$. One edge is $(E,D)$, the edge connecting these two vertices in the original Floquet code. This edge is of type $D$—the defect edge basis—and will not be included in the super-plaquette. Of the remaining four, two are of type $B$; these two will contribute to one of the shrink plaquettes, and will therefore be removed from the lattice during Steps 6 and 7. This leaves two edges incident to edge $(E,B)$, which by construction have to be a different type to $B$ or $D$, and these are the two edges that are included in the super-plaquette. Therefore the super-plaquette alternates before and after each edge in $C$. Furthermore, the cycle continues to alternate around the merge plaquette, by construction of the original code. In this way we see that the entire super-plaquette is formed from alternating edges of two bases. It can also be seen that this algorithm results in an updated lattice with the required properties of tri-valence and 3-face-colourability. We explain this in detail in Appendix B.
There are a few remarks we can note about this strategy. First of all, we can verify that no additional logical degrees of freedom have been created. To see this, note that two physical qubits are removed for the defect. After applying the strategy in the example of Fig. 2, four merge plaquettes are combined into a single super-plaquette, but two shrink plaquettes are split into two weight-2 plaquettes, meaning that the number of plaquettes has only decreased by one. We have also removed seven edges (one red, three blue, three green) and added four new edges (two blue, two green). Overall this means that we have removed one edge of each type from the graph, so after each sub-round the number of stabilisers formed by the edges of that sub-round also decreases by one. As a result, the size of the instantaneous stabiliser group after each sub-round has also decreased by two, meaning that the number of logical qubits—the number of physical qubits minus the size of the instantaneous stabiliser group during each sub-round—remains unchanged. As we show in Appendix B, this strategy is in fact optimal in minimising the number of physical qubits, plaquettes and connections removed while meeting the requirements for constructing a 3-face-colourable, tri-valent lattice.
@@
We can also note that while there are new edges added to the graph, these edges are between vertices which were adjacent in the original graph. This means that all the new two-body measurements are performed between qubits which are already connected on the quantum processor, implying that no additional hardware connectivity is required. This algorithm can also be run using the same measurement schedule as the original code, as the new measurements and plaquettes follow the same rules of tri-valence and 3-face-colourability as the original Floquet code.
@@
We can see that for multiple defects this algorithm can be implemented iteratively, by choosing a defect edge for each defective qubit. If multiple defects are close to each other, this can mean the super-plaquettes are themselves merged together to create even larger plaquettes. An example of two super-plaquettes merged together is shown in Fig. 3. Note that in this example the two defect edges selected are in the same basis, however this does not need to be true in general, and in certain cases it can be beneficial to choose different bases.
@@ 4.1 Percolation threshold
We estimate the probability of successfully preserving the logical observables using Monte Carlo sampling. First, we generate instances of the honeycomb code lattice for target code distances ranging from $d=3$ to $d=7$. We then select fabrication defect probabilities up to $14\%$. For each distance and defect rate, we generate up to 10,000 samples of defective qubits, and used the method described in Sec. 3 to adjust the code for them. We then estimate the probability that, when the algorithm was applied, the resulting graph did not percolate. Finally, we estimate a percolation threshold using second-order approximation methods, as described in Chapter 4 of Ref. [37]. Since this is only a property of the graph and the algorithm described above, the resulting percolating probabilities should apply to all Floquet codes defined on this lattice, not just the honeycomb code.
@@ 4.3 Pauli error threshold
Figure 7: Estimating a Pauli error threshold at $p_{\text{defect}}=6\%$. (a) Rescaled data $x=(p-p_{0})d^{1/\nu_{0}}$ is plotted against average logical error probability, along with a quadratic best fit. (b) Average logical error probability for various Pauli noise parameters, with estimated Pauli threshold (and error) in red.
@@
Despite the small average effective distances for $p_{\textrm{defect}}\geq 10\%$, we find clear evidence for sub-threshold behaviour at $p_{\textrm{defect}}=10\%$ (see Appendix F). The evidence for $p_{\textrm{defect}}>10\%$ is less clear, although our fits indicate a positive threshold.

### v2

@@ 1 Introduction
Floquet codes are a relatively new and exciting area in quantum error correction (QEC). From a theoretical perspective, they have a mathematical structure which is distinct from other families of codes such as stabiliser and subsystem codes, which lends them some interesting and unique properties [1]. In addition, from a more practical perspective, Floquet codes benefit from considerably reduced connectivity requirements in hardware, compared to the surface or colour codes, while maintaining a reasonably high threshold under Pauli noise and a low tera-quop footprint for certain noise models [2, 3, 4]. Finally, from a computational perspective, they provide simple constructions for implementing logical operations from the Clifford group [5].
@@
We numerically benchmark the performance of our method by implementing circuits for a planar honeycomb code memory experiment (specifically considering the model of Gidney et al. [3]) in Stim [26], a fast Clifford circuit simulator, and correcting for Pauli errors on the adapted code using PyMatching, a matching decoder [27]. We find considerable robustness to fabrication defects using our method: we determine a defect threshold of approximately $13\%$, and numerically demonstrate that Pauli error thresholds exist for defect probabilities close to this threshold. Hence, the region of parameter space in which the code remains fault-tolerant is large and comparable to that of the surface code [18]. The reduction in Pauli error threshold with defect probability is also quantitatively similar to previous findings for the surface code [20]. This is quite surprising, as the honeycomb code achieves this with more restrictive constraints on the connectivity of the device.
We further introduce two methods which take advantage of the underlying hardware. First, we propose even better methods for accommodating defects in hardware in which extra connectivity is present, such as on square lattices [28, 29, 30]. And secondly, we propose a method for better accommodating defective auxiliary qubits in lattices in which those are present—or, more generally, defective connections between data qubits—such as the heavy hex lattice [31].
The rest of this paper is laid out as follows. In Sec. 2 we provide some background on the honeycomb code (as a prototypical example from the family of Floquet codes we consider) and fabrication defects. In Sec. 3 we describe our strategy for removing defective qubits from the code lattice. Results from simulations are described in Sec. 4. We propose some improvements that can be made depending on the hardware of the quantum processor in Sec. 5. Finally, we conclude with some future research directions in Sec. 6.
@@ 2 Background
A Floquet code is defined by a series of check measurements and a sequence in which they are measured***For our purposes, we do not distinguish between finite or infinite-period dynamical codes, as other authors do [32, 5]; they will both be, in principle, amenable to our techniques.. Many Floquet codes are definable on the lattice of a colour code [24]. We focus here on the honeycomb code [1, 2], however this example can be easily generalised.
@@ 2.1 Honeycomb code
The honeycomb code is a Floquet code defined on a honeycomb latticeThis can be generalised to any lattice that is 3-face-colourable and tri-valent, i.e., where each vertex is adjacent to three edges.. Each vertex represents a physical qubit, and each edge represents a two-qubit Pauli measurement between neighbouring qubits. An example lattice is presented in Fig. 1, with the plaquettes coloured red, blue and green in such a way that no neighbouring plaquettes share the same colour (the lattice is 3-face-colourable). Edges are also assigned colours according to the plaquettes they connect; for example a red edge connects a pair of red plaquettes. A key benefit that this code has over other error-correcting codes, such as the surface code, is the low connectivity requirements of the code: each qubit only needs to be connected to three nearest neighbours (the lattice is tri-valent), and each measurement is only between pairs of adjacent qubits. This makes Floquet codes appealing for architectures with low connectivity, such as IBM’s heavy hex lattice [31], as well as devices with native two-body measurements such as Majorana devices [33, 4].
@@
The single-qubit measurements anti-commute with plaquettes along the boundary. The result of this is that some stabilisers are non-deterministic at times and therefore cannot detect errors at those times. However, this does not seem to significantly affect the code’s performance when compared to the periodic honeycomb code [3].
We form detectors—a comparison of two stabiliser measurements that can determine whether certain types of errors have occurred between the measurements—from two measurements of plaquette operators. For example, after the sequence $B\to R\to B$, we measure all bulk green plaquettes twice (once after $B\to R$, and a second time after $R\to B$); therefore detectors can be formed for each green plaquette by comparing the two measured values. An $X$ or $Z$ error occurring on a qubit in the plaquette’s support between the first and last of these three measurement sub-rounds will trigger the detector. Meanwhile, measurement results from detectors are deterministic in the presence of zero noise. A decoder can use these results to identify and correct errors when noise is present. Boundary plaquettes are measured non-deterministically when the single-qubit edge measurements precede the two-qubit edge measurements. However, they are re-measured deterministically (assuming no error has occurred) when the two-qubit edge measurements precede the single-qubit measurements. For example, a green boundary plaquette in Fig. 1 is first measured non-deterministically when the blue edge measurements immediately precede the red edge measurements ($B\to R$). It is then measured deterministically for a second time after the sequence $R\to B$, at which point we can form a detector comparing these two measurement outcomes.
@@ 2.2 Fabrication defects
Fabrication defects have been studied extensively in the context of surface codes, and can be accommodated by merging adjacent stabilisers together into “super-stabilisers” in order to remove the defective qubit(s) [18, 19, 20, 21]. These super-stabilisers are measured by introducing “gauge operators” around a defect, such that the super-stabiliser measurements can be formed by taking products of gauge measurements. This is similar to how edge measurements combine to form plaquette measurements in the honeycomb code, with the difference being that there is no dependency on the order in which gauge measurements are performed. However, translating this procedure to the honeycomb code is non-trivial: the adaptation of the lattice and the measurement schedule of the Floquet code must be performed carefully to make sure that detectors can still be formed near the defect locations. Meanwhile, although each sub-round of the honeycomb code can be mapped to an instance of a surface code [1], there are several impediments and disadvantages to directly adapting the surface code strategy, including that gauge operators of the surface code would not always correspond to edge measurements in the honeycomb code, and that the surface code mapped to from the honeycomb code evolves with each sub-round. Meanwhile, the resulting honeycomb code strategy may not be optimal, and the applicability of this strategy beyond the honeycomb code may also be limited. Hence, we must go beyond a simple adaptation of existing strategies to accommodate fabrication defects in Floquet codes.
Aasen et al. have investigated ways of accommodating fabrication defects on honeycomb and 4.8.8 Floquet codes [23]. They propose several strategies, two of which involve removing all qubits in a plaquette adjacent to a defect. An important consideration for such methods is whether the device has extra connectivity available for the formation of new plaquettes, which is required for some methods proposed [23]. Respecting the constraints of hardware while removing the smallest number of qubits and edges from the lattice possible is important for maintaining a high percolation threshold, and presumably a high Pauli noise threshold under finite fabrication defect ratesOne of the methods proposed in Ref. [23], while requiring extra connectivity, only necessitates the removal of a single qubit, which could be useful for honeycomb code implementations. However, the properties of the resulting graph and the updated measurement schedule make this strategy potentially hard to generalise to other Floquet codes (including the honeycomb code with boundaries) or to multiple defects in the lattice..
@@ 3 Strategy for removing defective qubits from the lattice
Here we outline the strategy for removing defective qubits from the honeycomb code’s underlying lattice. The main idea is to create “super-plaquettes” which avoid using the defective qubits, thereby removing them from the lattice. We do this by merging some plaquettes in the original lattice together, while other existing plaquettes are replaced by two-body plaquettes. We will do all this while retaining the 3-face-colourability and tri-valence of the graph, for ease of defining the updated code and ease of generalisation to other Floquet codes. To achieve this, we need to remove a pair of qubits along an edge, as described below. There is some subtlety around choosing the extra removed qubit (or qubits if there is more than one defect) in such a way that maximises the distance of the resulting code—we shall refer to this as the “edge choice”, since choosing a qubit to remove is the same as choosing an edge adjacent to the given fabrication defect. We also do not allow for extra connectivity between qubits in the updated Floquet code. In Appendix A, we describe how improved results can be obtained in the presence of extra connections in the underlying hardware.
@@ 3.1 The super-plaquette algorithm
The final result of this process is shown in Fig. 2(c). Note that the defective qubit and its neighbour are no longer connected to any other qubits, nor do they contribute to any plaquettes in the code. This means that they can now be removed from the code.
@@
The resulting super-plaquette is of the same type as the merge plaquettes: it is formed from a cycle of edges of alternating type not equal to $D$ (alternating $YY$ and $ZZ$ measurements for the example in Fig. 2). This follows from the fact that the cycle of edges that wraps around the super-plaquette is formed of edges from the set $C$ and edges from merge plaquettes, none of which have basis equal to the merge plaquette type. To see that the cycle of bases is alternating, consider the edges adjacent to $(E,B)\in C$ in Fig. 2(b). There are five edges adjacent to $E$. One edge is $(E,D)$, the edge connecting these two vertices in the original Floquet code. This edge is of type $D$—the defect edge basis—and will not be included in the super-plaquette. Of the remaining four, two are of type $B$; these two will contribute to one of the shrink plaquettes, and will therefore be removed from the lattice during Steps 6 and 7. This leaves two edges incident to edge $(E,B)$, which by construction have to be a different type to $B$ or $D$, and these are the two edges that are included in the super-plaquette. Therefore the super-plaquette alternates before and after each edge in $C$. Furthermore, the cycle continues to alternate around the merge plaquette, by construction of the original code. In this way we see that the entire super-plaquette is formed from alternating edges of two bases.
3.1.1 Remarks on super-plaquette algorithm
It can be seen that the algorithm introduced above results in an updated lattice with the required properties of tri-valence and 3-face-colourability. It does so while (at least for a defect within the bulk of the lattice) removing the smallest number of qubits, edges and plaquettes from the lattice possible. In Appendix B, we prove the following:
Theorem 1.
The algorithm above produces a tri-valent, 3-face-colourable lattice. It does so while removing the minimal number of qubits, edges and plaquettes from the lattice (assuming a lattice without boundaries).
@@ 3.1.1 Remarks on super-plaquette algorithm
We also note that while there are new edges added to the graph, these edges are between vertices which were adjacent in the original graph. This means that all the new two-body measurements are performed between qubits which are already connected on the quantum processor, implying that no additional hardware connectivity is required. This algorithm can also be run using the same measurement schedule as the original code, as the new measurements and plaquettes follow the same rules of tri-valence and 3-face-colourability as the original Floquet code.
An important consideration for maintaining a high threshold under fabrication defect noise is the number of connections removed from the lattice: that is, some qubits are connected by edges in the original lattice but not in the updated lattice. These removed connections are shown in white in Fig. 2(c). For uniform lattices (i.e., honeycomb, 4.8.8 or 4.6.12), the super-plaquette algorithm (again, applied in the bulk), results in the minimal removal of such connections without requiring additional connectivity in hardware. In Appendix B, we prove the following:
Theorem 2.
Consider a defective qubit in the bulk of a Floquet code defined on a honeycomb, 4.8.8 or 4.6.12 lattice (without boundary); i.e., a uniform tiling of the plane. Under the constraint that we cannot form an edge between any two qubits not connected by an edge in the original lattice, the algorithm above (with a certain choice of defect edge), minimises the number of removed edges from the original lattice.
There are a few further remarks we can note about this strategy. First of all, we can verify that no additional logical degrees of freedom have been created. To see this, note that two physical qubits are removed for the defect. After applying the strategy in the example of Fig. 2, four merge plaquettes are combined into a single super-plaquette, but two shrink plaquettes are split into two weight-2 plaquettes, meaning that the number of plaquettes has only decreased by one. We have also removed seven edges (one red, three blue, three green) and added four new edges (two blue, two green). Overall this means that we have removed one edge of each type from the graph, so after each sub-round the number of stabilisers formed by the edges of that sub-round also decreases by one. As a result, the size of the instantaneous stabiliser group after each sub-round has decreased by two generators, meaning that the number of logical qubits—the number of physical qubits minus the size of the instantaneous stabiliser group during each sub-round—remains unchanged.
@@
We can also see that for multiple defects this algorithm can be implemented iteratively, by choosing a defect edge for each defective qubit. If multiple defects are close to each other, this can mean the super-plaquettes are themselves merged together to create even larger plaquettes. An example of two super-plaquettes merged together is shown in Fig. 3. Note that in this example the two defect edges selected are in the same basis, however this does not need to be true in general, and in certain cases it can be beneficial to choose different bases (see Appendix D).
@@ 4.1 Percolation threshold
We estimate the probability of successfully preserving the logical observables using Monte Carlo sampling. First, we generate instances of the honeycomb code lattice for target code distances ranging from $d=3$ to $d=7$. We then select fabrication defect probabilities up to $14\%$. For each distance and defect rate, we generate up to 10,000 lattice samples, and use the method described in Sec. 3 to adjust the code for them. We then estimate the probability that, when the algorithm is applied, the resulting graph does not percolate. Finally, we estimate a percolation threshold using second-order approximation methods, as described in Chapter 4 of Ref. [37]. Since this is only a property of the graph and the algorithm described above, the resulting percolating probabilities should apply to all Floquet codes defined on this lattice, not just the honeycomb code.
@@ 4.3 Pauli error threshold
Figure 7: Estimating a Pauli error threshold at $p_{\text{defect}}=6\%$. (a) Rescaled data $x=(p-p_{0})d^{1/\nu_{0}}$ is plotted against average logical error probability, along with a quadratic best fit. Error bars indicate standard error of the mean for the 1,000 $p_{L}$ values sampled; the errors associated with each $p_{L}$ estimate, taken from 100,000 quantum memory experiment samples, is comparatively negligible. (b) Average logical error probability for various Pauli noise parameters, with estimated Pauli threshold (and error) in red.
@@
Despite the small average effective distances for $p_{\textrm{defect}}\geq 10\%$, we find clear evidence for sub-threshold behaviour at $p_{\textrm{defect}}=10\%$ (see Appendix F). The evidence for $p_{\textrm{defect}}>10\%$ is less clear, although our fits, and the increase in effective distance with target distance for $p_{\textrm{defect}}>10\%$ (Fig. 6), indicate a positive threshold.